OpenStudy (anonymous):

What is the vertex of this quadratic function? f(x) = 2x2 + 4x − 5

OpenStudy (anonymous):

first coordinate of the vertex is \(-\frac{b}{2a}\) which in your case is \(-\frac{4}{2\times 2}=-1\)

OpenStudy (anonymous):

second coordinate of the vertex is what you get when you replace \(x\) by the first coordinate of the vertex

OpenStudy (anonymous):

thank you! @satellite73